OpenStudy (anonymous):

What is the product of 2the square root of the quantity 10 times k. • 5kthe square root of 2.? Simplify if possible. @Nurali

OpenStudy (anonymous):

@zepdrix

zepdrix (zepdrix):

Like this? :o\[\Large 2\sqrt{10k}\cdot5k\sqrt{2}\]

OpenStudy (anonymous):

yes :o

zepdrix (zepdrix):

So we have a big bundle of multiplication here. Since multiplication is `commutative`, we can multiply in any order. Let's start by moving the 5k to the front and then multiply the 5 and 2 together.\[\Large 2\cdot5k\sqrt{10k}\cdot\sqrt{2} \qquad=\qquad 10k\sqrt{10k}\cdot\sqrt2\]

OpenStudy (anonymous):

@zepdrix ? o.o

zepdrix (zepdrix):

We can break down the first square root term like this, \[\Large 10k\sqrt{k}\sqrt{10}\cdot\sqrt2\] Let's break down the 10 a little bit,\[\Large 10k\sqrt{k}\sqrt{5\cdot2}\cdot\sqrt2 =10k\sqrt{k}\sqrt{5}\color{royalblue}{\sqrt{2}\cdot\sqrt2}\] Confused by any of those steps? Do you understand what this blue portion gives us when we multiply it out?

OpenStudy (anonymous):

I don't understand the second part.

zepdrix (zepdrix):

I rewrote 10 as 5times2. Is that the part that was confusing? :o

zepdrix (zepdrix):

Oh oh maybe this would be easier...

zepdrix (zepdrix):

\[\Large 10k\sqrt{10k}\cdot\sqrt2 =10k\sqrt{10k\cdot2}\]Combining the roots by bringing the 2 in. Now we can multiply the 2 and 10,\[\Large 10k\sqrt{20k}\]Does this way make a little bit more sense maybe? :O

zepdrix (zepdrix):

From here, there is a factor of 20 that is a perfect square. So we'll factor our 20 into 4x5.\[\Large 10k\sqrt{4\cdot5k}\]Taking the square root of 4 gives us? :o

OpenStudy (anonymous):

2 x 5k :o ?

zepdrix (zepdrix):

\[\Large 10k\cdot2\sqrt{5k}\] Good good. So as a final step it looks like you can multiply the 10 and 2 :O

OpenStudy (anonymous):

20k! :) Thanks.

zepdrix (zepdrix):

Yay good job \c:/ \[\Large 20k\sqrt{5k}\]